OpenStudy (anonymous):

Solve the question below. Let f(x)=x^2+3 and g(x)=x+2/x. Find (f o g)(2). When you get your answer, could you please show me the steps? Thank you.

OpenStudy (anonymous):

Here are my possible answers. A. 9 B. 9/7 C. 7 D. 14

OpenStudy (mathmale):

Lala: so that I know where you're coming from, please share what you know of that operator o as in (f o g)(2).

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

@mathmale f(g(x))

OpenStudy (anonymous):

For composite functions, I teach my student to write the OUTSIDE FUNCTION except where there is an X put a set of parenthesis. \[\left( \right)^{2}+3\] The next step is to put the INSIDE FUNCTION inside the parenthesis \[\left( \frac{ x+2 }{x } \right)^{2}+3\] This is your new COMPOSITE FUNCTION. Now substitute 2 in for X and you will get your answer.

OpenStudy (anonymous):

\[( )^2 +3 \] \[(x+2/x)^2 + 3\] \[16/4 + 3\] \[4+3=7\]

OpenStudy (anonymous):

Yes, that is your answer.

OpenStudy (anonymous):

@hullsnipe Thank you! Your steps were so easy to follow. I have another question coming up, could you also help me with that one?